Use the following data to calculate (a) the size of the labour force and (b) the official unemployment rate: total population, 500; population under 15 years of age or institutionalized, 120; not in labour force, 150; unemployed, 23; part-time workers looking for full-time jobs, 10.
1
Expert's answer
2021-09-13T12:45:13-0400

a) Size of labour force= Total population"-" Population Under 16 years of age or Institutionalized"-" Not in labour force.


"=500-120-150=230"


So total labour force is 230


b) Calculating unemployment rate by using Unemployed=23, and total Labour force=230.

That is, "\frac {Unemployed} {Total labour force} ×100"


"=\frac{23}{230}\times 100 =10\%"


So the unemployment rate is 10%